Please show me how to set up an equation for 4x -y = 9
x= 2.
I'm not sure how to start this: Am I multiplying or subtracting? What are the steps?
4(2)-y=9

Answer by rfer(16322)

good start, keep going
8-y=9
-y=9-8
-y=1
y=-1